# Bayesian decomposition using Besov priors

### v1

@@ 1 Introduction
Inverse problems concern the estimation of unknown parameters of a system given indirect measurements. Such problems are typically ill-posed, meaning the computation of the solution is heavily sensitive to measurement limitations, measurement noise, and imprecise models. Stable solutions are often obtained through regularization techniques, which imposes penalties on the solution, e.g., periodicity, smoothness, nonsmoothness, or sparsity with respect to an expansion representation. In many inverse imaging problems the unknown parameters naturally consist of multiple components, e.g., image decomposition of cartoon and texture parts; see, for example, [1, 2, 3, 4, 5]. Taking the multiple components into account when modeling such inverse problems enables the use of a specific regularizer on each component to achieve a balanced solution that avoids overfitting to a subset of the components.
@@
where $\mathbf{y}\in\mathbb{R}^{m}$ is the data, $\mathbf{f}\in\mathbb{R}^{n}$ is the unknown which is decomposed into the two components $\mathbf{g},\mathbf{h}\in\mathbb{R}^{n}$, $\mathbf{A}\in\mathbb{R}^{m\times n}$ is the linear forward operator, and $\mathbf{\epsilon}\in\mathbb{R}^{m}$ is additive noise. We assume that the component $\mathbf{g}$ is piecewise constant and the component $\mathbf{h}$ is smooth. These assumptions are reasonable in imaging applications such as anomaly detection in image deblurring [6, 7], atmospheric emission tomography [8], magnetoencephalography [9], and X-ray computed tomography [10].
Many variational methods have been proposed for signal or image decomposition [11, 12, 13, 14], but all of these methods lack the ability to quantify uncertainties in the decomposition results with respect to noisy data. In this work, we take the Bayesian approach to solve inverse problems as described in [15], that is, the inverse problem is formulated in a probabilistic sense in terms of random variables and their associated distributions. Regularization is now imposed via the prior probability distribution on the unknown, while model inaccuracies and measurement noise are accounted for in the likelihood function. The solution to a Bayesian inverse problem, called the posterior distribution, is the conditional distribution of the unknown given an observation of the data. Assuming that the components $\mathbf{g}$, $\mathbf{h}$ and the noise $\mathbf{\epsilon}$ are mutually independent, the posterior distribution associated with the inverse problem (1) is given by Bayes’ theorem [15] as
@@
The formulation of prior distributions in Bayesian inverse problems is crucial for obtaining accurate estimates of the unknown and ensuring that its assumed properties are properly enforced. In this work, we focus on two classes of priors: smoothness priors for representing the component $\mathbf{h}$, and priors that promote the piecewise constant structure for representing the component $\mathbf{g}$. One way of modeling smoothness priors is to construct prior distributions that impose smoothness regularity in a specific function space. These models include Gaussian priors that are realized in Sobolev spaces [16, 17] and Besov priors that are realized in Besov spaces [18, 19]. These models offer a consistent representation of smoothness and provide explicit control over it through the choice of prior parameters. The modeling of priors that promote piecewise constants is an active area of research, with several approaches available such as heavy-tailed priors [20, 21, 22, 23, 24], hierarchical shrinkage priors [25, 26, 27, 28, 29], machine-learning priors [30, 31, 32], and Haar-wavelet-based Besov priors [33, 34, 35, 36]. Each of these methods has its strengths and limitations and should be selected based on the specific constraints of the problem at hand. We focus primarily on Besov priors as they offer a flexible framework for modeling both smooth and piecewise constant structures, while also allowing efficient computation and interpretable parameter choices.
@@
The selection of parameters, which controls the relative contribution of each component’s prior information, is crucial for obtaining a balanced decomposition under the two-component model. In classical regularization methods, these parameters are often tuned manually based on visual quality, and there has been limited work on automatic parameter selection for decomposition, e.g., [8]. However, in the Bayesian framework, we can include these parameters as hyperparameters and estimate them jointly with the decomposition solution via the posterior. The main challenges lie in how to design priors and hyperpriors to ensure the components are well separated.
@@
In this paper, we propose two Bayesian methods to solve linear inverse problems, where the unknown consists of two components with different properties. In the first method, we introduce a novel way of combining two distinct Besov priors by formulating a two-Besov decomposition model. We utilize the Besov prior with a Haar wavelet basis to impose jump discontinuities and the Besov prior with a smooth wavelet to impose smoothness. The No-U-Turn sampler (NUTS)[37] was applied to sample the posterior. In the second method, we propose a new hierarchical posterior to incorporate the parameters that control the strengths of the priors as hyperparameters. In this method, we use a combination of a hierarchical Gaussian prior on the discrete derivative to impose sparsely located jumps and a Besov prior with a smooth wavelet to impose localized smoothness. To sample this hierarchical posterior, we introduce a Gibbs sampling scheme, where marginal distributions either have closed-form expressions or can be effectively sampled using a Randomize-Then-Optimize (RTO) method [38, 36]. Both methods are tested on 1D and 2D deconvolution problems. The numerical results indicate that the decomposition method yields better reconstructions compared to methods using a single Besov prior. Furthermore, the hierarchical posterior can achieve automatic parameter selection for decomposition problems.
@@ 2.1 Besov priors and wavelets
Following [39], we let $\phi,\psi\in C^{r}(\mathbb{R})$ with $r\geq 1$ be compactly supported scaling and wavelet functions, respectively. There exist choices of such wavelets $\psi$, where the associated scaling function $\phi$ generates a multiresolution analysis (MRA) of $L^{2}(\mathbb{R})$. An example is the Daubechies wavelet of order $M$, denoted by DB(M)-wavelet in the following, which is supported in $[0,2M-1]$ and belongs to $C^{\gamma M}(\mathbb{R})$ for $\gamma\approx 0.2075$ [40].
@@
where $j\in\mathbb{N}_{0}$, $k\in\mathbb{Z}^{d}$ and $\ell=(\ell_{1},\ell_{2},\ldots,\ell_{d})\in\mathcal{L}_{0}$. Note that $\ell=(0,0,\ldots,0)$ corresponds to the scaling function, while each $\ell\in\mathcal{L}_{1}$ corresponds to the $12d-$ 1 wavelet functions. The collection $\left\{\psi_{0,k,(0,0,\ldots,0)}\},\;k\in\mathbb{Z}^{d}\right\}\cup\left\{\psi_{j,k,\ell}\mid j\in\mathbb{N}_{0},\;k\in\mathbb{Z}^{d},\;\ell\in\mathcal{L}_{1}\right\}$ forms an orthonormal basis of $L^{2}(\mathbb{{R}}^{d})$ [41].
@@
Set the index set as $\mathcal{I}_{\psi}=\{(0,0,0,\ldots,0)\}\cup\{j\in\mathbb{N}_{0},k\in\mathcal{K}_{j},\ell\in\mathcal{L}_{1}\}$ then the collection $\{\psi^{\text{per}}_{j,k,\ell},(j,k,\ell)\in\mathcal{I}_{\psi}\}$ is an orthonormal basis of $L^{2}(\mathbb{T}^{d})$ [42]. The 1-periodic orthonormal wavelet basis can, under conditions on its smoothness, be used to define a class of smoothness spaces known as Besov spaces [43].
@@
[…]
@@
The random wavelet expansion in (11) almost surely takes values in the space $B_{p,p}^{t}(\mathbb{T}^{d})$ if and only if $t<s-\frac{d}{p}$ [18] and generates a Besov probability measure defined by a pushforward measure $\mathbb{Q}\circ T^{-1}$ [19]. Formally, this gives rise to a Besov prior with Lebesgue density given by
@@ 2.2 Discrete Besov priors
We will use the fast wavelet transform [44] implemented in the Pywavelets module [45] for computing the wavelet transform with $\mathbf{W}_{\psi}$. The continuous Besov norm can be approximated by a discrete version as
@@ 2.3 The posterior for Bayesian decomposition
The solution to the Bayesian decomposition problem is given by the posterior distribution, which combines the prior and likelihood. According to Bayes’ theorem [15], the posterior is proportional to the product of the likelihood and the prior, i.e.,
@@
When the prior strength parameters $\lambda_{g}$ and $\lambda_{h}$ are fixed, we use the No-U-Turn sampler (NUTS) [37] to sample from the posterior distribution in (19). NUTS is a Hamiltonian Markov chain Monte Carlo method with automatic adaptive step size control, which usually makes it efficient and eliminates the need for manual tuning of the sampler. The negative log posterior of (19) is almost everywhere differentiable for all choices of prior parameters $s_{g},s_{h}>0$ and $p_{g},p_{h}\in[1,\infty)$, which makes NUTS applicable in this setting.
In decomposition problems, the results strongly depend on the strength parameters $\lambda_{g}$ and $\lambda_{h}$, which define the balance between two components. The choice of the strength parameters is crucial and challenging. Under Bayesian statistics, we can introduce a hierarchical model to incorporate both parameters as unknown random variables, called hyperparameters, in the posterior [^1] The model is derived in A. For numerical experiments, see Figures 6 and 7.. Unfortunately, even when the two priors are constructed from structurally different wavelets, the resulting hierarchical model suffers from an identifiability problem. In general, two wavelet bases have high mutual coherence defined as the maximum absolute value of the inner product between two wavelet functions from different bases [46]. In particular, we see a large coherence between scaling functions and other low-frequency wavelets from the different bases. This high mutual coherence leads to a lack of identifiability in the posterior distribution. When the strength parameters are considered as random variables, it becomes difficult to distinguish the contributions between two Besov priors, as they can represent similar features in the unknown. This lack of identifiability can result in a multimodal posterior, as different parameter choices may cause different priors to dominate the reconstruction. Sampling from multimodal posteriors is a well-known challenge for Gibbs sampling and Markov chain Monte Carlo methods in general, and addressing it is beyond the scope of this paper. The identifiability issues are numerically illustrated in Section 4.
@@ 3 Hierarchical Gaussian-Besov decomposition method
where $\otimes$ denotes the Kronecker product. Referring to the work in [27], we consider a diagonal covariance $\mathbf{\Lambda}$ and assume that the diagonal entries are i.i.d. random variables following the inverse gamma distribution $\mathcal{IG}(\alpha_{1},\beta_{1})$. This hierarchical Gaussian prior imposes sparsity on the gradient of $\mathbf{g}$. For the smooth component $\mathbf{h}$, we continue to use a discrete Besov prior $\mathbf{h}\sim T_{p_{h},\psi_{h}}^{s_{h}}$ with $p_{h}=2$ and a smooth wavelet $\psi_{h}$. Furthermore, we set the prior strength parameter $\lambda_{h}$ as a hyperparameter following the gamma distribution $\mathcal{G}(\alpha_{2},\beta_{2})$, which is conjugate to the Gaussian distribution. Combining with the Gaussian likelihood, we form the joint posterior of the components and hyperparameters as
@@
To sample the marginal distributions $\pi_{1}$ and $\pi_{2}$, we notice that they are Gaussian and can be sampled efficiently using the Randomize-Then-Optimize (RTO) method [38, 36]. RTO draws samples from a Gaussian distribution by solving the following random perturbed linear least squares problems
@@
where $\bm{\eta}^{*}$ and $\bm{\zeta}^{*}$ are realizations of $\bm{\eta}$ and $\bm{\zeta}$, respectively. We solve the linear least squares problems using a CGLS algorithm [47]. To sample the covariance matrix $\mathbf{\Lambda}$ we notice that $\pi_{3}$ is an independent product of
@@ 4.1 One dimensional deconvolution: Two-Besov prior method
The ground truth signal $\mathbf{f}_{\text{truth}}$ is constructed as a linear combination of Haar wavelets and DB(8) wavelets, discretized on a uniform grid of 512 points over the interval $[0,1)$ with a truncation parameter set to $J=9$. The convolution operator $A$ is discretized on the same grid, resulting in a discrete forward operator $\mathbf{A}\in\mathbb{R}^{512\times 512}$. This operator is implemented as a circulant matrix corresponding to the Gaussian convolution kernel in (28) truncated at ±$\pm 3\sigma_{\text{ker}}$ with $\sigma_{\text{ker}}=0.01$. The degraded data $\mathbf{y}^{*}$ is generated by adding $10\%$ Gaussian noise $\epsilon$, i.e., $\frac{\|\mathbf{\epsilon}\|_{2}}{\|\mathbf{A}\mathbf{f}_{\text{true}\|_{2}}}=0.1$, on $\mathbf{A}\mathbf{f}_{\text{true}}$. Figure 1 displays the ground truth signal $\mathbf{f}_{\text{true}}$, its components $\mathbf{g}$ and $\mathbf{h}$, and the noisy observation $\mathbf{y}^{*}$.
@@
For all experiments in this test, inference is performed using NUTS to generate 15 0001500015\,00015 000 total samples, of which the first $5000$ are discarded as burn-in. The remaining 10 0001000010\,00010 000 samples are used to compute the posterior mean estimate, the 95% credible interval (CI), and the CI width. To assess sampling quality, we compute the effective sample size (ESS) over the entire grid and report the minimum, median, and maximum values. Additionally, we evaluate the autocorrelation function (ACF) at a representative grid point to examine sample correlation. Reconstruction quality is measured using the relative error between the posterior mean and the ground truth $\mathbf{f}_{\text{true}}$.

### v2

@@ 1 Introduction
Inverse problems concern the estimation of unknown parameters of a system given indirect measurements. Such problems are typically ill-posed, meaning the computation of the solution is heavily sensitive to measurement limitations, measurement noise, and imprecise models. Stable solutions are often obtained through regularization techniques, which imposes penalties on the solution, e.g., periodicity, smoothness, nonsmoothness, or sparsity with respect to an expansion representation. In many inverse imaging problems the unknown parameters naturally consist of multiple components, e.g., image decomposition of cartoon and texture parts; see, for example, (?, ?, ?, ?, ?). Taking the multiple components into account when modeling such inverse problems enables the use of a specific regularizer on each component to achieve a balanced solution that avoids overfitting to a subset of the components.
@@
where $\mathbf{y}\in\mathbb{R}^{m}$ is the data, $\mathbf{f}\in\mathbb{R}^{n}$ is the unknown which is decomposed into the two components $\mathbf{g},\mathbf{h}\in\mathbb{R}^{n}$, $\mathbf{A}\in\mathbb{R}^{m\times n}$ is the linear forward operator, and $\mathbf{\epsilon}\in\mathbb{R}^{m}$ is additive noise. We assume that the component $\mathbf{g}$ is piecewise constant and the component $\mathbf{h}$ is smooth. These assumptions are reasonable in imaging applications such as anomaly detection in image deblurring (?, ?), atmospheric emission tomography (?), magnetoencephalography (?), and X-ray computed tomography (?).
Many variational methods have been proposed for signal or image decomposition (?, ?, ?, ?), but all of these methods lack the ability to quantify uncertainties in the decomposition results with respect to noisy data. In this work, we take the Bayesian approach to solve inverse problems as described in (?), that is, the inverse problem is formulated in a probabilistic sense in terms of random variables and their associated distributions. Regularization is now imposed via the prior probability distribution on the unknown, while model inaccuracies and measurement noise are accounted for in the likelihood function. The solution to a Bayesian inverse problem, called the posterior distribution, is the conditional distribution of the unknown given an observation of the data. Assuming that the components $\mathbf{g}$, $\mathbf{h}$ and the noise $\mathbf{\epsilon}$ are mutually independent, the posterior distribution associated with the inverse problem (1) is given by Bayes’ theorem (?) as
@@
The formulation of prior distributions in Bayesian inverse problems is crucial for obtaining accurate estimates of the unknown and ensuring that its assumed properties are properly enforced. In this work, we focus on two classes of priors: smoothness priors for representing the component $\mathbf{h}$, and priors that promote the piecewise constant structure for representing the component $\mathbf{g}$. One way of modeling smoothness priors is to construct prior distributions that impose smoothness regularity in a specific function space. These models include Gaussian priors that are realized in Sobolev spaces (?, ?) and Besov priors that are realized in Besov spaces (?, ?). These models offer a consistent representation of smoothness and provide explicit control over it through the choice of prior parameters. The modeling of priors that promote piecewise constants is an active area of research, with several approaches available such as heavy-tailed priors (?, ?, ?, ?, ?), hierarchical shrinkage priors (?, ?, ?, ?, ?), machine-learning priors (?, ?, ?), and Haar-wavelet-based Besov priors (?, ?, ?, ?). Each of these methods has its strengths and limitations and should be selected based on the specific constraints of the problem at hand. We focus primarily on Besov priors as they offer a flexible framework for modeling both smooth and piecewise constant structures, while also allowing efficient computation and interpretable parameter choices.
@@
The selection of parameters, which controls the relative contribution of each component’s prior information, is crucial for obtaining a balanced decomposition under the two-component model. In classical regularization methods, these parameters are often tuned manually based on visual quality, and there has been limited work on automatic parameter selection for decomposition, e.g., (?). However, in the Bayesian framework, we can include these parameters as hyperparameters and estimate them jointly with the decomposition solution via the posterior. The main challenges lie in how to design priors and hyperpriors to ensure the components are well separated.
@@
In this paper, we propose two Bayesian methods to solve linear inverse problems, where the unknown consists of two components with different properties. In the first method, we introduce a novel way of combining two distinct Besov priors by formulating a two-Besov decomposition model. We utilize the Besov prior with a Haar wavelet basis to impose jump discontinuities and the Besov prior with a smooth wavelet to impose smoothness. The No-U-Turn sampler (NUTS)(?) was applied to sample the posterior. In the second method, we propose a new hierarchical posterior to incorporate the parameters that control the strengths of the priors as hyperparameters. In this method, we use a combination of a hierarchical Gaussian prior on the discrete derivative to impose sparsely located jumps and a Besov prior with a smooth wavelet to impose localized smoothness. To sample this hierarchical posterior, we introduce a Gibbs sampling scheme, where marginal distributions either have closed-form expressions or can be effectively sampled using a Randomize-Then-Optimize (RTO) method (?, ?). Both methods are tested on 1D and 2D deconvolution problems. The numerical results indicate that the decomposition method yields better reconstructions compared to methods using a single Besov prior. Furthermore, the hierarchical posterior can achieve automatic parameter selection for decomposition problems.
@@ 2.1 Besov priors and wavelets
Following (?), we let $\phi,\psi\in C^{r}(\mathbb{R})$ with $r\geq 1$ be compactly supported scaling and wavelet functions, respectively. There exist choices of such wavelets $\psi$, where the associated scaling function $\phi$ generates a multiresolution analysis (MRA) of $L^{2}(\mathbb{R})$. An example is the Daubechies wavelet of order $M$, denoted by DB(M)-wavelet in the following, which is supported in [0,2M1][0,2M-1] and belongs to $C^{\gamma M}(\mathbb{R})$ for $\gamma\approx 0.2075$ (?).
@@
where $j\in\mathbb{N}_{0}$, $k\in\mathbb{Z}^{d}$ and $\ell=(\ell_{1},\ell_{2},\ldots,\ell_{d})\in\mathcal{L}_{0}$. Note that $\ell=(0,0,\ldots,0)$ corresponds to the scaling function, while each $\ell\in\mathcal{L}_{1}$ corresponds to the 2d12d-1 wavelet functions. The collection $\left\{\psi_{0,k,(0,0,\ldots,0)}\},\;k\in\mathbb{Z}^{d}\right\}\cup\left\{\psi_{j,k,\ell}\mid j\in\mathbb{N}_{0},\;k\in\mathbb{Z}^{d},\;\ell\in\mathcal{L}_{1}\right\}$ forms an orthonormal basis of $L^{2}(\mathbb{{R}}^{d})$ (?).
@@
Set the index set as $\mathcal{I}_{\psi}=\{(0,0,0,\ldots,0)\}\cup\{j\in\mathbb{N}_{0},k\in\mathcal{K}_{j},\ell\in\mathcal{L}_{1}\}$ then the collection $\{\psi^{\text{per}}_{j,k,\ell},(j,k,\ell)\in\mathcal{I}_{\psi}\}$ is an orthonormal basis of $L^{2}(\mathbb{T}^{d})$ (?). The 1-periodic orthonormal wavelet basis can, under conditions on its smoothness, be used to define a class of smoothness spaces known as Besov spaces (?).
@@
$\pi_{\xi}(x)=\frac{1}{c_{p}}\exp\left(-|x|^{p}\right),\;c_{p}=\int_{\mathbb{R}}\exp\left(-|x|^{p}\right)dx,\;x\in\mathbb{R}.$ (9)
@@
The random wavelet expansion in (11) almost surely takes values in the space $B_{p,p}^{t}(\mathbb{T}^{d})$ if and only if $t<s-\frac{d}{p}$ (?) and generates a Besov probability measure defined by a pushforward measure $\mathbb{Q}\circ T^{-1}$ (?). Formally, this gives rise to a Besov prior with Lebesgue density given by
@@ 2.2 Discrete Besov priors
We will use the fast wavelet transform (?) implemented in the Pywavelets module (?) for computing the wavelet transform with $\mathbf{W}_{\psi}$. The continuous Besov norm can be approximated by a discrete version as
@@ 2.3 The posterior for Bayesian decomposition
The solution to the Bayesian decomposition problem is given by the posterior distribution, which combines the prior and likelihood. According to Bayes’ theorem (?), the posterior is proportional to the product of the likelihood and the prior, i.e.,
@@
When the prior strength parameters $\lambda_{g}$ and $\lambda_{h}$ are fixed, we use the No-U-Turn sampler (NUTS) (?) to sample from the posterior distribution in (19). NUTS is a Hamiltonian Markov chain Monte Carlo method with automatic adaptive step size control, which usually makes it efficient and eliminates the need for manual tuning of the sampler. The negative log posterior of (19) is almost everywhere differentiable for all choices of prior parameters $s_{g},s_{h}>0$ and $p_{g},p_{h}\in[1,\infty)$, which makes NUTS applicable in this setting.
In decomposition problems, the results strongly depend on the strength parameters $\lambda_{g}$ and $\lambda_{h}$, which define the balance between two components. The choice of the strength parameters is crucial and challenging. Under Bayesian statistics, we can introduce a hierarchical model to incorporate both parameters as unknown random variables, called hyperparameters, in the posterior [^1] The model is derived in A. For numerical experiments, see Figures 6 and 7.. Unfortunately, even when the two priors are constructed from structurally different wavelets, the resulting hierarchical model suffers from an identifiability problem. In general, two wavelet bases have high mutual coherence defined as the maximum absolute value of the inner product between two wavelet functions from different bases (?). In particular, we see a large coherence between scaling functions and other low-frequency wavelets from the different bases. This high mutual coherence leads to a lack of identifiability in the posterior distribution. When the strength parameters are considered as random variables, it becomes difficult to distinguish the contributions between two Besov priors, as they can represent similar features in the unknown. This lack of identifiability can result in a multimodal posterior, as different parameter choices may cause different priors to dominate the reconstruction. Sampling from multimodal posteriors is a well-known challenge for Gibbs sampling and Markov chain Monte Carlo methods in general, and addressing it is beyond the scope of this paper. The identifiability issues are numerically illustrated in Section 4.
@@ 3 Hierarchical Gaussian-Besov decomposition method
where $\otimes$ denotes the Kronecker product. Referring to the work in (?), we consider a diagonal covariance $\mathbf{\Lambda}$ and assume that the diagonal entries are i.i.d. random variables following the inverse gamma distribution $\mathcal{IG}(\alpha_{1},\beta_{1})$. This hierarchical Gaussian prior imposes sparsity on the gradient of $\mathbf{g}$. For the smooth component $\mathbf{h}$, we continue to use a discrete Besov prior $\mathbf{h}\sim T_{p_{h},\psi_{h}}^{s_{h}}$ with $p_{h}=2$ and a smooth wavelet $\psi_{h}$. Furthermore, we set the prior strength parameter $\lambda_{h}$ as a hyperparameter following the gamma distribution $\mathcal{G}(\alpha_{2},\beta_{2})$, which is conjugate to the Gaussian distribution. Combining with the Gaussian likelihood, we form the joint posterior of the components and hyperparameters as
@@
To sample the marginal distributions $\pi_{1}$ and $\pi_{2}$, we notice that they are Gaussian and can be sampled efficiently using the Randomize-Then-Optimize (RTO) method (?, ?). RTO draws samples from a Gaussian distribution by solving the following random perturbed linear least squares problems
@@
where $\bm{\eta}^{*}$ and $\bm{\zeta}^{*}$ are realizations of $\bm{\eta}$ and $\bm{\zeta}$, respectively. We solve the linear least squares problems using a CGLS algorithm (?). To sample the covariance matrix $\mathbf{\Lambda}$ we notice that $\pi_{3}$ is an independent product of
@@ 4.1 One dimensional deconvolution: Two-Besov prior method
The ground truth signal $\mathbf{f}_{\text{truth}}$ is constructed as a linear combination of Haar wavelets and DB(8) wavelets, discretized on a uniform grid of 512 points over the interval $[0,1)$ with a truncation parameter set to $J=9$. The convolution operator $A$ is discretized on the same grid, resulting in a discrete forward operator $\mathbf{A}\in\mathbb{R}^{512\times 512}$. This operator is implemented as a circulant matrix corresponding to the Gaussian convolution kernel in (28) truncated at $\pm 3\sigma_{\text{ker}}$ with $\sigma_{\text{ker}}=0.01$. The degraded data $\mathbf{y}^{*}$ is generated by adding $10\%$ Gaussian noise $\epsilon$, i.e., $\frac{\|\mathbf{\epsilon}\|_{2}}{\|\mathbf{A}\mathbf{f}_{\text{true}\|_{2}}}=0.1$, on $\mathbf{A}\mathbf{f}_{\text{true}}$. Figure 1 displays the ground truth signal $\mathbf{f}_{\text{true}}$, its components $\mathbf{g}$ and $\mathbf{h}$, and the noisy observation $\mathbf{y}^{*}$.
@@
For all experiments in this test, inference is performed using NUTS to generate 15 00015\,000 total samples, of which the first $5000$ are discarded as burn-in. The remaining 10 00010\,000 samples are used to compute the posterior mean estimate, the 95% credible interval (CI), and the CI width. To assess sampling quality, we compute the effective sample size (ESS) over the entire grid and report the minimum, median, and maximum values. Additionally, we evaluate the autocorrelation function (ACF) at a representative grid point to examine sample correlation. Reconstruction quality is measured using the relative error between the posterior mean and the ground truth $\mathbf{f}_{\text{true}}$.